OpenStudy (anonymous):

\[\LARGE f'(x)=x^2+3x \;\; ,\;\; f(1)=5\;\;\;\;\Longrightarrow f(2)=?\] I need some hints

OpenStudy (anonymous):

f(x) can be obtained by integration........then use f(1) to find value of c, then find f(2)

OpenStudy (anonymous):

ya suroj ur correctt

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[\LARGE f(x)=\frac13 x^3 +\frac32 x^2 \;\;\;\;??\]

OpenStudy (anonymous):

you left +C at the end

OpenStudy (anonymous):

you need to add +C at end because it is integration

OpenStudy (anonymous):

indefinite integration

OpenStudy (anonymous):

ok but... I need to firnd f(2) then how??

OpenStudy (anonymous):

on the second condition when x=1, y=5, you see , put those values and get c

OpenStudy (anonymous):

\[\LARGE \int\limits x^2+3x \;dx = \frac13 x^3+\frac32x^2+C\]

OpenStudy (anonymous):

put the c back on f(x) and get complete f(x) equation and then find f(2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

how can I know what the value of C is ??

OpenStudy (anonymous):

I see it you told me before... let me try, I'll see what I can do :D

OpenStudy (anonymous):

see second condition f(1) is 5 than mean when x=1 , y=5

OpenStudy (anonymous):

see f(x) is what you get by integration

OpenStudy (anonymous):

because you integrate f'(x)

OpenStudy (anonymous):

did you catch it?

OpenStudy (anonymous):

is C=19/6 ??

OpenStudy (anonymous):

I'm still dizzy ! O_O

OpenStudy (anonymous):

yes you got it

OpenStudy (anonymous):

now put that C back in f(x)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ahh yeah... now I'll do it thank you very much ;)

OpenStudy (anonymous):

no problem lol

OpenStudy (anonymous):

\[f \left( x \right) = x ^{2} + 3x\] \[f \left( 1 \right) = 1^{2} + 3 \times 1 = 5\] \[f \left( 2 \right) = 2^{2} + 3 \times 2 = 10\] simple.

OpenStudy (anonymous):

@kreshnik another simple way to do something.

OpenStudy (anonymous):

@JoshDavoll WRONG... the answer IS \[\frac{71}{6}\] !! Thanks for the help !

OpenStudy (anonymous):

hahaha... obviously , you mus be 10-12 years ... LOL